Coronado Furniture Company started construction of a combination office and warehouse building for its own use at an estimated cost of $7,500,000 on January 1, 2020. Coronado expected to complete the building by December 31, 2020. Coronado has the following debt obligations outstanding during the construction period.

Construction loan-12% interest, payable semiannually, issued December 31, 2019 $3,000,000
Short-term loan-10% interest, payable monthly, and principal payable at maturity on May 30, 2021 2,100,000
Long-term loan-11% interest, payable on January 1 of each year. Principal payable on January 1, 2024 1,500,000

Assume that Coronado completed the office and warehouse building on December 31, 2020, as planned at a total cost of $7,800,000, and the weighted-average amount of accumulated expenditures was $5,400,000. Compute the avoidable interest on this project.

Compute the depreciation expense for the year ended December 31, 2021. Coronado elected to depreciate the building on a straight-line basis and determined that the asset has a useful life of 30 years and a salvage value of $450,000.
Answer #1
Calculation of weighted average interest rate:
Type of loan Principal Interest Rate Interest
Short Term Loan-10% Interest                                                       2,100,000 10%          210,000
Long Term Loan-11% Interest                                                       1,500,000 11%          165,000
Total                                                       3,600,000          375,000
Weighted Average Interest Rate= Total Interest / Total Principal
=375000/3600000
=10.42%
Available Interest Calculation:
Particulars Principal Balance of Weighted Average Accumulated Expenditure Interest Rate Available Interest
Interest on Loan specifically for Contruction                                                       3,000,000 12%          360,000
Interest on remaining loan upto weighted average expenditure total using weighted average interest rate                                                       2,400,000 10.42%          250,080
Total                                                       5,400,000          610,080
Available Interest $ 610,080
Calculation of Depreciation Expense:
Total Cost =Cost Incurred + Interest Capitalised
=7800000+610080
=8410080
Depreciation Expense =(Total Cost-Estimated Salvage Value)/Estimated Useful Life
=(8410080-450000)/30
Depreciation Expense =265336